OpenStudy (anonymous):

-24/55 divided 39/40

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

-24/55 times 40/39

OpenStudy (anonymous):

-960/2145 = -192/429 = -64/143

OpenStudy (anonymous):

how did you get -960/2145

OpenStudy (anonymous):

-24/55 * 40/39: -24*40 = 960 / 55*39 = 2145.

OpenStudy (anonymous):

okay and thank you very much